Solvetheequationsbelowandhence,showthesolutionsoftheseequationsontheArganddiagram:(i)z4−i=0(ii)z3=−√3−i(iii)z5=−18i(iv)z=16i1+i18(v)z=1+i√3+i16


1
Expert's answer
2021-05-07T10:09:43-0400

(i)"z4-i=0"

"4(x+iy)-i=0\Rightarrow 4x+(4y-1)i=0\Rightarrow x=0 \text{ and }y=\dfrac{1}{4}"


(ii) "z3=-\sqrt{3}-i"

"3(x+iy)+\sqrt{3}+i=0\Rightarrow (3x+\sqrt{3})+i(y+1)=0\Rightarrow x=\dfrac{-1}{\sqrt{3}}, y=-1"


(iii)"z5=-18i"

"5(x+iy)+18i=0\Rightarrow 5x+i(y+18)=0\Rightarrow x=0,y=-18"


(iv) "z=16i+18i"

"x+iy=34i\Rightarrow x=0,y=34"


(v) "z=1+i\sqrt{3}+16i"

"x+iy=1+i(\sqrt{3}+16)\Rightarrow x=1,y=\sqrt{3}+16"